What current is required to produce 200 liter of hydrogen gas measured at S.T.P from electrolysis of water in 0.5 hours


1
Expert's answer
2021-07-01T05:31:03-0400

W = I x t

W is the weight of the substance

I is current

t is time for which current flows

I = W / t

n (H2) = V / 22.4 = 200 / 22.4 = 8.93 mol

n = m / M

M (H2) = 2 g/mol

m (H2) = n x M = 8.9 x 2 = 17.9 g

I = W / t = 17.9 / (30 x 60) = 0.01 A